OpenStudy (anonymous):

If y varies directly as x, and y = 25 as x = 5, find y for the x-value 7.

OpenStudy (anonymous):

do the same?

OpenStudy (anonymous):

now here y=kx just find k bu putting the values of x and y

OpenStudy (anonymous):

here 25=k*5 or k=5 thus y=5x now if x=7 then y=5*7=35

OpenStudy (anonymous):

for direct variation y= kx for inverse variation y=k/x

OpenStudy (anonymous):

You're the man dude. Freaking thanks!